OpenStudy (anonymous):

Write the equation of the line that passes through each pair of points in slope-intercept form. (-1,7),(4,-3)

OpenStudy (s):

First you have to find a slope(m): m=y2-y2/x2-x1 m=(-3-7)/4-(-1) m=-10/5 m=-2 now you know that y-y1=m(x-x1) pick up any point you have (-1;7) or (4;-3) for example, i will take (-1;7) y-7=-2(x-(-1) y-7=-2(x+1) y-7=-2x-2 y=-2x-2+7 y=-2x+7 --> this is your line that passes through your two given points